OpenStudy (anonymous):

divide and simply (2h)^9/(2h)^11

OpenStudy (anonymous):

Okay for this problem we use the rule of exponents: a^m/a^n=a^m-n

OpenStudy (anonymous):

Therefore , 2h^9/2h^11=2h^9-11=2h^-2=1/2h^2